I am lost on this equation. Can someone help?


Which of the ordered pairs (-3,3), (9,0), (-3,-3), (-23,8)
are solutions for the equation x + 4y = 9?
Thank you, Sher

Answer by funmath(2933)

Hi Sher,
Isn't this your last week? Finish Strong girl!!!
Which of the ordered pairs (-3,3), (9,0), (-3,-3), (-23,8)
are solutions for the equation x + 4y = 9?
When (x,y)=(-3,3)
-3+4(3)=9 ?
-3+12=9
9=9 (-3,3) is a solution.
:
When (x,y)=(9,0)
9+4(0)=9
9+0=9
9=9 (9,0) is also a solution.
:
When (x,y)=(-3,-3)
-3+4(-3)=9
-3-12=9
-15=9 (-3,-3) is not a solution.
:
When (x,y)=(-23,8)
-23+4(8)=9
-23+32=9
9=9 (-23,8) is a solution.
:
Happy Calculating!!!